OpenStudy (anonymous):

35 less than seven of a certain number is fifty more than 10 times of that number. what is the number

OpenStudy (firejay5):

@slawhorne What did you get as your answer

OpenStudy (anonymous):

none, could not even get started

OpenStudy (firejay5):

Look in the sentence for keys terms to (add, subtract, multiply, or divide)

OpenStudy (anonymous):

ok. what I got is 35-7x=10x+50?

OpenStudy (firejay5):

What grade are you in and what math subject?

OpenStudy (anonymous):

9th, algebra 1

OpenStudy (firejay5):

Really, I never did this in Algebra I What is this thing called that you're having to do

OpenStudy (anonymous):

this the problem gave the teacher asked to try to do, because tomorrow we will have quiz on these types of problems. what is it called-brain teaser

OpenStudy (firejay5):

Here let me give you another example, so you know how to work it

OpenStudy (anonymous):

ok

OpenStudy (firejay5):

Try this: If three hens lay three eggs in three days, how many eggs does a (statistical) hen lay in one day?

OpenStudy (anonymous):

<1

OpenStudy (firejay5):

yes it's less than 1, but what number exactly (Hint: Fraction)

OpenStudy (anonymous):

1/3

OpenStudy (firejay5):

bam!!!! You Nailed it :D

OpenStudy (firejay5):

One third. (Note: 3 hens = 3 eggs / 3 days → 3 hens = (3 / 3) (eggs / days) → 1 hen = (1 / 3) (egg / days))

OpenStudy (anonymous):

thank you, I have stop now and finish my hw, good night

OpenStudy (firejay5):

Another answer would be: Zero or one (it's hard to lay a third of an egg). On your question: Here's what I got: 7x=85+10x -3x=85 x=-85/3

OpenStudy (firejay5):

make sure to give me a medal lol! :D @slawhorne